OpenStudy (anonymous):

A certain capacitor in a circuit has a capacitive reactance of 36.8 Ω when the frequency is 136 Hz. What capacitive reactance does the capacitor have at a frequency of 20.3 kHz?

OpenStudy (anonymous):

I keep coming up with the wrong answer I use Xc = 1/(2*pi*F*C) rearrange and solve for C: C = 1/(2*pi*F*Xc) F = frequency Xc = capacitave reactance Then Xc = 1/(2*pi*F*C) solve for Xc using C found above and same frequency

OpenStudy (anonymous):

didnt convert my 20.3 to kHz. solved.

OpenStudy (kropot72):

The reactance of the capacitor is inversely proportional to frequency. Reactance at 1 Hz = 36.8 * 136 ohms Reactance at 20.3 kHz = \[\frac{36.8\times 136}{20.3\times 10^{3}}=you\ can\ calculate\]